Suppose the market demand function is given by D(PD) = 60—PD and the supply is S(P) = 2PS.

a) Find the market equilibrium without taxes where PD = PS

b) Suppose now the supplier is required to a per unit tax of t = 3. Draw a graph to to show the change in the supply curve. Find the equilibrium quantity and prices, q*,PS*,and PD*. How much of the per unit tax is passed on to the consumer?

c) Suppose now that the demander is required to a per unit tax of t = 3. Draw a graph to show the change in the demand curve Find the equilibrium quantity and prices, q*, PS*, and PD* .Compare your answers in parts b) and c).

d) Calculate the change in consumer’s surplus, change in producer’s surplus, and dead weight loss as a result of the policy change. Illustrate them on a graph.


1
Expert's answer
2021-11-24T12:09:33-0500

a) The market equilibrium without taxes where PD = PS and D(P) = S(P) is:

60—P = 2P,

3P = 60,

P = 20,

Q = 2×20 = 40 units.